OpenStudy (klepez):

simplify -11 multiplyed by the square root of 112

OpenStudy (klepez):

here are the choices \[-44\sqrt{7}\]

OpenStudy (klepez):

\[-176\sqrt{7}\]

OpenStudy (klepez):

\[-27\sqrt{7}\]

OpenStudy (klepez):

and \[4\sqrt{7}\]

OpenStudy (anonymous):

\[-11\sqrt{112}\]\[-11\sqrt{16}\sqrt{7}\]\[-44\sqrt{7}\]

OpenStudy (klepez):

thank you

OpenStudy (anonymous):

ur welcome :)